Build a line which will represent the graph of the equation y–2.5x=c, if it’s known that it passes through the point K (2, –3).
Answer by fractalier(6550)

If the line passes thru (2, -3), let us plug in and find c...
-3 - 2.5(2) = c
-3 - 5 = c
c = -8
so that
y - 2.5x = -8
or
y = 2.5x - 8